# Hierarchical Clustering With Confidence

### v1

@@ 5 Simulation and real data analysis
We turn now to empirical evaluations of the methodology developed in the previous sections. We begin, in Section 5.1, by studying the effect of different levels of randomization on the quality of a clustering. In Section 5.2, we evaluate the validity and power of our proposed conditional p-value. Section 5.3 assesses the effectiveness of proposed procedure for choosing $K$. Finally, Section 5.4 applies our method in a real-world scenario.
@@ 5.1 How does randomization affect clustering quality?
A natural question is how the introduced randomization impacts the quality of the resulting clustering. To investigate this, we evaluate clustering performance across different values of $\tau$ in Algorithm 1 using the following two metrics:
  1. (i)
  2. (ii)
    $\text{ARI}=\dfrac{\sum_{k,k^{\prime}}\binom{n_{k,k^{\prime}}}{2}-\left[\sum_{k}\binom{a_{k}}{2}\sum_{j}\binom{b_{k^{\prime}}}{2}\right]\Big/\binom{n}{2}}{\dfrac{1}{2}\left[\sum_{k}\binom{a_{k}}{2}+\sum_{k^{\prime}}\binom{b_{k^{\prime}}}{2}\right]-\left[\sum_{k}\binom{a_{k}}{2}\sum_{j}\binom{b_{k^{\prime}}}{2}\right]\Big/\binom{n}{2}}$,
We generate 500 synthetic datasets, each consisting of $n=30$ observations in $p=2$ dimensions, drawn from $\mathcal{N}(\mu,\sigma^{2}I_{p})$ with variance $\sigma^{2}=1$. The true number of clusters in this setting is $K=2$. Following the simulation setup in (Yun and Barber, 2023), the data observation are divided into two clusters
@@
where $\delta>0$ is a parameter controlling the signal strength. Each cluster contains $n/2=15$ observations.
First, we apply randomized complete linkage hierarchical clustering with the true number of clusters, $K=2$, fixing $\delta=6$, and compare the clustering performance for $\tau\in\{0,0.025,0.05,0.1,0.25,0.5,1,5\}$. Note that $\tau=0$ corresponds to the usual (deterministic) agglomerative clustering algorithm, depicted as RC(0), while the nonzero values of $\tau$, in ascending order, correspond to randomization levels 11-77, denoted RC(1) through RL(7).
Figure 2: Comparison of clustering quality metrics under varying levels of randomization $\tau$. (Left): Boxplots of the ratio between within-cluster sum of squares (WCSS) and total sum of squares (TSS), showing how cluster compactness changes with $\tau$. (Right): Boxplots of the Adjusted Rand Index (ARI), measuring agreement with the true clustering, which declines as $\tau$ increases.
The results, shown in Figure 2, indicate that for small values of $\tau$, the randomized clustering yields results that are comparable to the deterministic agglomerative clustering in terms of both compactness of the clustering structure and label recovery. Consistent with expectations, as $\tau$ increases, the external randomness begins to dominate the merging process, resulting in less stable clusters and a clear decline in clustering quality, as reflected by higher WCSS/TSS and lower ARI. For example, a sudden decline in clustering quality is observed for RC(4), corresponding to $\tau=0.25$. These patterns are consistent across all three metrics used to evaluate clustering performance.
We also compare the performance of the randomized hierarchical clustering algorithm under varying levels of signal strength $\delta$, while fixing the randomization level at $\tau=0.10$ (i.e., implementing RC(3)). Figure 3 shows that the randomized hierarchical clustering algorithm with $\tau=0.10$ achieves performance very similar to the deterministic version across a wide range of signal strengths $\delta$. When the clusters are well separated (larger $\delta$), the randomized method attains nearly perfect recovery of the true clustering, as reflected by low WCSS/TSS ratios and high ARI values. Overall, these results suggest that a small amount of randomization does not degrade clustering quality, even as the signal strength varies. The benefits of the added randomization, however, become evident in the next sections, where we demonstrate the performance of our testing procedures for validating the clusters discovered by Algorithm 1.
@@ 5.2 Evaluating p-value validity and power
We now empirically evaluate the validity and power of the proposed conditional p-value, examining its null distribution, Type I error control, and detection power under different alternatives.
@@
We conduct two main experiments: (i) we first show that our test yields valid p-values for any degree of randomization; (ii) we then compare the ECDF of the p-values obtained at randomization level $\tau=0.10$ (i.e., RC(3)) with the adjusted chi-squared test from Gao et al. (2024), which uses the plug-in $\sigma^{2}$ estimate $\widehat{\sigma}^{2}=\frac{1}{(n-1)p}\sum_{i=1}^{n}\|X_{i}-\bar{X}\|_{2}^{2}$, and with the adjusted F test from Yun and Barber (2023). We include the naive p-values to illustrate that they suffer from the double-dipping issue and consequently yield invalid tests across all scenarios in both experiments.
To obtain the ECDF, we perform 2000 independent trials, generating the rows of the data matrix $X$ in each trial as independent draws:
$X_{i}\sim\mathcal{N}(\mu_{i},\sigma^{2}I_{p}),\quad i\in\{1,2,\ldots,n\}.$
For each $i$, the mean vectors are set to $\mu_{i}=0_{p}$, corresponding to the null hypothesis being true. We fix $\sigma=1$, $n=30$, and $p=10$. Alignment with the $y=x$ line in the ECDF plot indicates the validity of the p-values. To evaluate Type I error control, we adopt the same data generation process as before. For each repetition, we perform $200$ trials to estimate the empirical Type I error rate, and we repeat this process $100$ times to construct the boxplots.
In the first experiment, we perform complete-linkage hierarchical clustering with RC(1)-RC(7) until the data are partitioned into $K=2$ and $K=3$ clusters, respectively, and compute p-values for the resulting clusters. As shown in Figure 4, the proposed method produces p-values that are uniformly distributed under the null hypothesis and successfully controls the Type I error rate, at all levels of randomization. In contrast, the naive method yields excessively small p-values with a severely inflated Type I error rate.
@@
In the second experiment, we compare the proposed p-values obtained with the proposed randomized method with $\tau=0.1$ (i.e., RC(3)) against the p-values from the naive method and the p-values obtained from the selective inferential approaches in (Gao et al., 2024) and (Yun and Barber, 2023). Using the same data-generating model as before with $n=30$, $p=10$, $\sigma=1$, we perform our clustering algorithm until the data are partitioned into $K=2$ and $K=3$ clusters, respectively, under four linkage criteria: complete, average, single, and minimax. For each setting, we perform $2000$ repetitions and evaluate the ECDF of the resulting p-values to assess validity under the null. The software implementations of the methods proposed by (Gao et al., 2024) and (Yun and Barber, 2023) do not support testing clusters obtained using the minimax linkage criterion. Figure 5 shows that, across all settings, the ECDFs of the p-values from RC(3) align closely with the reference uniform distribution, indicating validity. As expected, the results are similar to the two existing selective inferential methods whenever these approaches are available for a given linkage criterion.
Figure 5: Comparison of the ECDFs of p-values obtained under the null hypothesis. The proposed method with $\tau=0.1$ is compared against (Gao et al., 2024) and (Yun and Barber, 2023), across different linkage criteria and cluster numbers $K\in\{2,3\}$.
@@
To evaluate power, we consider two scenarios in which the true number of clusters is $K\in\{2,3\}$. For $K=2$, data are generated according to (16) with $n=30$ and $\sigma=1$. For $K=3$, data are generated from three equidistant clusters
We consider nine evenly spaced values of $\delta\in[1,10]$ and simulate $2000$ datasets for each $\delta$. For each dataset, we conduct the hypothesis test in (4) at a significance level of $\alpha=0.05$.
We compute the probability of rejecting the alternative as a function of effect size, which we define as
$\text{ES}:=d(M;\mu),$
where $d(M;\mu)$ is the dissimilarity between the true cluster means $\mu$, based on the linkage function employed in the clustering algorithm. We note that $\text{ES}=0$ under the null for any linkage, as $d(M;\mu)=0$.
To estimate the rejection probability as a function of effect size, we bin the simulated outcomes into $10$ equally spaced intervals based on their effect sizes. Within each bin, we calculate the proportion of rejections to obtain the estimated power. Confidence intervals are constructed for each bin using a normal approximation to the binomial distribution. Figure 6 summarizes the empirical power of the proposed randomized method with $\tau=0.1$ (i.e., RC(3)) across different linkage functions and true cluster numbers $K\in\{2,3\}$, and compares it against the two other existing selective inference methods.
Figure 6: Empirical power curves as a function of effect size for the proposed randomized method with $\tau=0.1$, compared with the two selective inference approaches under varying choice of linkage functions (complete, single, average and minimax) and true number of clusters $K\in\{2,3\}$.
@@ 5.3 Evaluating our procedure for choosing the number of clusters
We next evaluate the empirical performance of our estimation procedure for determining the number of clusters, as presented in Algorithm 2, using the conditional p-values developed in Section 3. We first demonstrate that the sequential testing procedure achieves valid family-wise error rate (FWER) control under the global null, and then compare its ability to estimate the number of clusters with that of the widely used gap statistic proposed in (Tibshirani et al., 2001) across a wide range of settings.
In our experiments, we employ an exponentially decaying $\alpha$-sequence of the form $\alpha_{j}\propto\exp{(-0.5j)}$ for $i=1,\dots,n-1$, normalized such that the total testing budget sums to $\alpha=0.05$. We set the cutoffs for the minimum testable cluster size and for deciding between a more liberal versus conservative choice of $\alpha$ in Algorithm 2 to $n_{\text{min}}=0.1\times n$ and $n^{*}=0.4\times n$, respectively. Throughout all experiments, the randomization level is fixed at $\tau=0.1$, corresponding to RC(3).
To assess FWER control, we generate data under the null model in (4) with $n=30$, $p=2$, and $\sigma=1$, so that no clustering structure is present. In this setting, selecting $\widehat{K}>1$ is equivalent to rejecting at least one true null hypothesis in the merge sequence. The empirical FWER is therefore the proportion of repetitions for which $\widehat{K}>1$. As shown in Table 1, across $2000$ independent repetitions, the proposed procedure, under various randomization levels, keeps this fraction below the target level $\alpha=0.05$. In contrast, Algorithm 2 with naive p-values (denoted Naive) fails to maintain this control.
@@
We next evaluate how accurately the method recovers the number of clusters when true clustering structure is present. We compare our approach with the results from using the gap statistic in Tibshirani et al. (2001), which selects $\widehat{K}$ by assessing whether the clusters obtained from the observed data are significantly tighter than those formed in reference datasets lacking structure. Specifically, for each candidate $K$, the gap statistic computes
$\text{Gap}(K)=\mathbb{E}_{\text{ref}}[\log W_{K}]-\log(W_{K}),$
where $W_{K}$ denotes the within‐cluster sum of squares. The selected number of clusters is then defined as:
$\widehat{K}_{\text{Gap}}=\min\{K:\text{Gap}(K)\geq\text{Gap}(K+1)-s_{K+1},\}$
where $s_{K}$ is the standard deviation adjustment term defined in (Tibshirani et al., 2001).
@@
Figure 7 demonstrates that our sequential testing procedure more consistently recovers the correct clustering structure than the gap statistic across all settings. As cluster separation increases, it nearly always identifies the true $K$. By contrast, the gap statistic remains overly conservative and frequently underestimates the true number of clusters, even for well-separated clusters.
We further evaluate the performance of Algorithm 2 under varying true numbers of clusters $K^{*}\in\{1,2,\dots,10\}$. For each value of $K^{*}$, we generate $100$ independent datasets with $n=200$ observations and $p=2$ dimensions. The $K^{*}$ cluster centers are placed evenly around a circle of radius $\delta=6$, and each data point is sampled from a spherical Gaussian distribution centered at its corresponding cluster. Cluster sizes are distributed as evenly as possible across groups. We use the same procedure as in the previous experiments to generate the $\alpha$-sequence, and set $n_{\min}=10$ and $n^{*}=40$.
@@ 5.4 Real data analysis
In this section, we analyze the penguin dataset (Horst et al., 2020). We focus on a subset of 107 female penguins observed during the years 2007–2008, using the three species Adelie, Chinstrap, and Gentoo as ground-truth labels. The clustering analysis is based solely on the two continuous measurements, bill length and flipper length, and does not use the three true labels.
@@
    To estimate the number of clusters $\widehat{K}$, we run $100$ independent repetitions of Algorithm 2 using complete linkage and a randomization level of $\tau=0.1$, and select the mode of the resulting distribution as our estimate $\widehat{K}$. We employ the same $\alpha$-sequence, $n_{\min}$, and $n^{*}$ as described in Section 5.3.
@@
    After selecting $\widehat{K}$, we repeatedly run the clustering procedure $500$ times to assess stability of the clustering results. The pairwise co-occurrence matrix is constructed as follows: for every pair of samples $(i,j)$, we count the number of times they are assigned to the same cluster across all runs and divide by the total number of runs. Sharper color contrasts between blocks indicate more stable clustering results.
Figure 9: (Left) Histogram of the estimated number of clusters $\widehat{K}$. The blue dashed line shows $\widehat{K}=2$ given by the gap statistics. (Right) Heatmap displaying the pairwise co-occurrence matrix across 500 independent trails using $\widehat{K}=2$ and $\tau=0.1$.
@@ 6 Conclusion
Validating clustering results is challenging because the same data cannot be used both to form the clusters and to test whether those clusters are genuine. In this paper, we address this challenge by developing tests that quantify the evidence against each merge in a hierarchical clustering algorithm. The resulting p-values are then used to determine when to stop the greedy merging process and, in turn, to estimate the number of clusters.
@@
At the core of our approach is a novel randomized hierarchical clustering method. Randomization gives researchers flexible control to retain more information for inference or validation, without compromising the quality of selection. In particular, the p-values produced by our testing procedure allow the same data to be used twice for both clustering and validation, resulting in more powerful tests while controlling the desired Type I error rate. Importantly, our empirical results indicate that this inference comes at little cost to clustering quality when the level of randomization is kept small. Such an approach avoids the typical tradeoff in data-splitting methods, where allocating too much data for selection can leave too little information for inference, while allocating too little data for selection can compromise clustering quality.
@@
Beyond the statistical advantages of using the same data for both clustering and validation, our proposed randomization scheme offers a practical strategy to construct tractable inference in hierarchical clustering. It can be applied with any linkage, avoiding both the derivation of linkage-specific tests and the use of Monte Carlo approximations for computing such tests. This gives our approach an advantage over existing selective inferential methods, which can only explicitly characterize the conditioning set for certain linkages and dissimilarities based on Euclidean distances, but must rely on costly sampling-based approximations for others.
@@
Finally, practical interpretation of the output from a hierarchical clustering method relies on selecting an appropriate dendrogram level or cut height in order to extract and assess the resulting cluster structure. To this end, the p-values developed in our paper—with the appropriate conditional Type I error guarantees—are used to construct a sequential testing procedure that selects the number of clusters with a rigorous probabilistic guarantee against overestimation.

### v2

@@ 5 Simulation and real data analysis
We turn now to empirical evaluations of the methodology developed in the previous sections. We begin, in Section 5.1, by studying the effect of different levels of randomization on the quality of a clustering. In Section 5.2, we evaluate the validity and power of our proposed conditional p-value. Section 5.3 assesses the effectiveness of proposed procedure for choosing $K$. Finally, Section 5.4 applies our method in a real-world scenario. Software implementations of our tests are available in Python at https://github.com/judywu4800/SI_HierarchicalClustering and in R at https://github.com/jacobbien/randhclust-project.
@@ 5.1 How does randomization affect clustering quality?
[…]
We generate 500 synthetic datasets, each consisting of $n=30$ observations in $p=2$ dimensions, drawn from $\mathcal{N}(\mu,\sigma^{2}I_{p})$ with variance $\sigma^{2}=1$. The true number of clusters in this setting is $K=2$. Following the simulation setup in [Yun and Barber, 2023], the data observation are divided into two clusters
@@
where $\delta>0$ controls the signal strength. Each cluster contains $15$ observations.
First, we apply complete-linkage randomized clustering with the true number of clusters, $K=2$ and $\delta=6$, and compare the clustering performance for $\tau^{*}\in\{0,0.025,0.05,0.1,0.25,0.5,1,5\}$. Note that $\tau^{*}=0$ corresponds to the traditional agglomerative clustering algorithm, depicted as RC(0), while the nonzero values of $\tau^{*}$, in ascending order, correspond to randomization levels 11-77, denoted RC(1) through RL(7).
Figure 2: Comparison of clustering quality metrics under varying levels of randomization $\tau^{*}$. (Left): Boxplots of the ratio between within-cluster sum of squares (WCSS) and total sum of squares (TSS), showing how cluster compactness changes with $\tau^{*}$. (Right): Boxplots of the Adjusted Rand Index (ARI), measuring agreement with the true clustering, which declines as $\tau^{*}$ increases.
The results, shown in Figure 2, indicate that for small values of $\tau^{*}$, the randomized clustering yields results that are comparable to the deterministic agglomerative clustering in terms of both compactness of the clustering structure and label recovery. Consistent with expectations, as $\tau^{*}$ increases, the external randomness begins to dominate the merging process, resulting in less stable clusters and a clear decline in clustering quality, as reflected by higher WCSS/TSS and lower ARI. For example, a sudden decline in clustering quality is observed for RC(4), corresponding to $\tau^{*}=0.25$. These patterns are consistent across all three metrics used to evaluate clustering performance.
[…]
@@ 5.2 Evaluating p-value validity and power
We now evaluate the validity and power of the proposed conditional p-value, focusing on its null distribution, Type I error control, and power under alternatives.
@@
[…]
To obtain the ECDF, we perform 2000 independent trials, generating the rows of the data matrix $X$ in each trial as independent draws: $X_{i}\sim\mathcal{N}(\mu_{i},\sigma^{2}I_{p}),\quad i\in\{1,2,\ldots,n\}$. For each $i$, the mean vectors are set to $\mu_{i}=0_{p}$, so that any selected null hypothesis is true. We fix $\sigma=1$, $n=30$, and $p=10$. Alignment with the $y=x$ line in the ECDF plot indicates the validity of the p-values. To evaluate Type I error control, we adopt the same data generation process as before. For each repetition, we perform $200$ trials to estimate the empirical Type I error rate, and we repeat this process $100$ times to construct the boxplots.
In the first experiment, we perform complete-linkage hierarchical clustering with RC(1)-RC(7) until the data are partitioned into $K=2$ clusters, respectively, and compute p-values for the resulting clusters. As shown in Figure 3, the proposed method produces p-values that are uniformly distributed under the null hypothesis and successfully controls the Type I error rate, at all levels of randomization. By contrast, the naive method yields excessively small p-values with a severely inflated Type I error rate.
@@
[…]
@@
To evaluate power, we generate data from $K=2$ true clusters according to (16) with $n=30$ and $\sigma=1$. We consider nine evenly spaced values of $\delta\in[1,10]$ and simulate $2000$ datasets for each $\delta$. For each dataset, we conduct the hypothesis test in (6) at a significance level of $\alpha=0.05$.
We compute the probability of rejecting the alternative as a function of effect size, which we define as $\text{ES}:=d(M;\mu)$, where $d(M;\mu)$ is the dissimilarity between the true cluster means $\mu\in\mathbb{R}^{n\times p}$, based on the linkage function employed in the clustering algorithm. We note that $\text{ES}=0$ under the null for any linkage, i.e. $d(M;\mu)=0$.
To estimate the rejection probability as a function of effect size, we bin the simulated outcomes into $10$ equally spaced intervals based on their effect sizes. Within each bin, we calculate the proportion of rejections to obtain the estimated power. Confidence intervals are constructed for each bin using a normal approximation to the binomial distribution. Figure 4 summarizes the empirical power of the proposed randomized method with $\tau^{*}=0.10$ (i.e., RC(3)) across different linkage functions and true cluster numbers $K=2$, and compares it against the two other existing selective inference methods.
Figure 4: Empirical power curves as a function of effect size for the proposed randomized method with $\tau^{*}=0.10$, compared with the two selective inference approaches under varying choice of linkage functions (complete, single, average and minimax) and true number of clusters $K=2$.
@@ 5.3 Evaluating our procedure for choosing the number of clusters
We next evaluate the empirical performance of our estimation procedure for determining the number of clusters, as presented in Algorithm 2, using the conditional p-values developed in Section 3. We first demonstrate that the sequential testing procedure achieves valid family-wise error rate (FWER) control under the global null, and then compare its ability to estimate the number of clusters with that of the widely used gap statistic proposed in Tibshirani et al. [2001] across a wide range of settings.
In our experiments, we employ an exponentially decaying $\alpha$-sequence of the form $\alpha_{j}\propto\exp{(-0.5j)}$ for $i=1,\dots,n-1$, normalized such that the total testing budget sums to $\alpha=0.05$. We set the cutoffs for the minimum testable cluster size and for deciding between a more liberal versus conservative choice of $\alpha$ in Algorithm 2 to $n_{\text{min}}=0.1\times n$ and $n^{*}=0.4\times n$, respectively. Throughout all experiments, the randomization level is fixed at $\tau^{*}=0.10$, corresponding to RC(3).
To assess FWER control, we generate data under the null model in (6) with $n=30$, $p=2$, and $\sigma=1$, so that no clustering structure is present. In this setting, selecting $\widehat{K}>1$ is equivalent to rejecting at least one true null hypothesis in the merge sequence. The empirical FWER is therefore the proportion of repetitions for which $\widehat{K}>1$. As shown in Table 1, across $2000$ independent repetitions, the proposed procedure, under various randomization levels, keeps this fraction below the target level $\alpha=0.05$. By contrast, Algorithm 2 with naive p-values (denoted Naive) fails to maintain this control.
@@
We next evaluate how accurately the method recovers the number of clusters when true clustering structure is present. We compare our approach with the results from using the gap statistic in Tibshirani et al. [2001], which selects $\widehat{K}$ by assessing whether the clusters obtained from the observed data are significantly tighter than those formed in reference datasets lacking structure. Specifically, for each candidate $K$, the gap statistic computes $\text{Gap}(K)=\mathbb{E}_{\text{ref}}[\log W_{K}]-\log(W_{K})$, where $W_{K}$ denotes the within‐cluster sum of squares. The selected number of clusters is then defined as $\widehat{K}_{\text{Gap}}=\min\{K:\text{Gap}(K)\geq\text{Gap}(K+1)-s_{K+1}\},$ where $s_{K}$ is the standard deviation adjustment term defined in [Tibshirani et al., 2001].
@@
[…]
@@ 5.4 Real data analysis
In this section, we analyze the penguin dataset [Horst et al., 2020] that was previously studied in Gao et al. [2024], Yun and Barber [2023]. We focus on a subset of 107 female penguins observed during the years 2007–2008, using the three species Adelie, Chinstrap, and Gentoo as ground-truth labels. The clustering analysis is based solely on the two continuous measurements, bill length and flipper length.
@@
    To estimate the number of clusters $\widehat{K}$, we run $100$ independent repetitions of Algorithm 2 using complete linkage and a randomization level of $\tau^{*}=0.10$, and select the mode of the resulting distribution as our estimate $\widehat{K}$. We employ the same $\alpha$-sequence, $n_{\min}$, and $n^{*}$ as described in Section 5.3.
@@
    After selecting $\widehat{K}$, we repeatedly run our randomized clustering procedure $500$ times to assess stability of the clustering results. The pairwise co-occurrence matrix is constructed as follows: for every pair of samples $(i,j)$, we count the number of times they are assigned to the same cluster across all runs and divide by the total number of runs. Sharper color contrasts between blocks indicate more stable clustering results.
Figure 6: (Left) Histogram of the estimated number of clusters $\widehat{K}$. The blue dashed line shows $\widehat{K}=2$ given by the gap statistics. (Right) Heatmap displaying the pairwise co-occurrence matrix across 500 independent trials using $\widehat{K}=2$ and $\tau^{*}=0.10$.
@@ 6 Conclusion
Validating clustering results is challenging because the same data cannot be used both to form the clusters and to test whether those clusters are genuine. In this paper, we address this challenge by developing tests that quantify the evidence against each merge in a hierarchical clustering algorithm. The resulting p-values are then used to determine when to stop the merging process and, in turn, to estimate the number of clusters.
@@
At the core of our approach is a novel randomized hierarchical clustering method. Randomization gives researchers flexible control to retain more information for inference or validation, without compromising the quality of selection. In particular, the p-values produced by our testing procedure allow the same data to be used twice for both clustering and validation, resulting in more powerful tests while controlling the desired Type I error rate.
@@
Beyond the statistical advantages of using the same data for both clustering and validation, our proposed randomization scheme offers a practical strategy to construct tractable inference in hierarchical clustering. It can be applied with any linkage without requiring the derivation of linkage-specific tests or the use of Monte Carlo approximations. This gives our approach an advantage over existing selective inferential methods, which can only explicitly characterize the conditioning set for certain linkages and dissimilarities based on Euclidean distances, but must rely on costly sampling-based approximations for others.
@@
Finally, interpreting hierarchical clustering output relies on selecting an appropriate dendrogram level or cut height in order to extract and assess the resulting cluster structure. To this end, the p-values developed in our paper—with the appropriate conditional Type I error guarantees—are used to construct a sequential testing procedure that selects the number of clusters with a rigorous probabilistic guarantee against overestimation.